# Aligning Human Motion Generation with Human Perceptions

### v1

@@ 1 Introduction
Human motion generation is an important emerging task [53] with wide-ranging applications, including augmented and virtual reality (AR/VR) [48, 46], human-robot interaction [29, 11], and digital humans [23, 47]. Achieving high realism in generated human motions is crucial, necessitating naturalness, smoothness, and plausibility. However, current generation methods still fall short of these goals, often producing subpar results. Meanwhile, designing appropriate evaluation metrics that accurately reflect these qualities remains a significant challenge. This complexity stems from the highly non-linear and articulated nature of human motion, which must adhere to physical and bio-mechanical constraints while also avoiding visual artifacts. Effective metrics would not only facilitate the objective comparison of generated results but also have the potential to enhance generation models by addressing their shortcomings.
@@
Existing evaluation metrics typically rely on error with pairing ground truth (GT) motion, simple heuristics, or on distribution distance with real motion manifold. The error-based metrics cannot fully reflect the performance because GT is only one reasonable possibility. The heuristics fall short in comprehensively representing motion quality. For instance, foot-ground contact metrics [32, 40] fail to penalize twisting arm motions that violate bio-mechanical constraints. It is also infeasible to manually define all the human motion rules in a handcrafted manner. Meanwhile, distribution distance metrics like Fréchet Inception Distance (FID) [15] do not operate on an instance level but rather assess overall distribution similarity. Consequently, they cannot identify implausible motions or provide direct supervision signals to guide the generation of higher-quality motions. Some studies [40, 41] also indicate that FID correlates poorly with user studies due to the misalignment between its distance measurement and human perception of motion quality. Consequently, existing automatic evaluation metrics cannot effectively reflect or replace subjective user studies, hindering objective evaluation and comparison.
@@
In light of this, we advocate the need for automatic evaluation aligned with human perceptions. Firstly, humans are the primary audience and interaction partners for motion generation, making their perception crucial for evaluating motion quality. Secondly, the human brain possesses specialized neural mechanisms for processing biological motion [3, 10] and is sensitive to even slightly unnatural motions [39, 34]. Therefore, we explore the possibility of directly learning perceptual evaluations from humans using a data-driven approach. This method could bridge the gap between objective metrics and subjective human judgments, providing a more accurate assessment of motion quality.
@@
First, we carefully curate a human perceptual evaluation dataset named MotionPercept, which contains $52590$ pairs of human preference annotations on generated motions. Next, we train a human motion critic model, MotionCritic, that learns motion quality ratings from the collected dataset. Our critic model significantly outperforms previous metrics in terms of alignment with human perceptions. Notably, it generalizes well across different data distributions. In addition to motion evaluation, we further propose to utilize the critic model as a direct supervision signal. We demonstrate that MotionCritic can be seamlessly integrated into the generation training pipeline, effectively improving motion generation quality by increasing alignment with human perceptions with few steps of finetuning.
@@ 2.1 Human Motion Generation
Human motion generation is a pivotal task in computer vision, computer graphics, and artificial intelligence, aiming to produce natural and realistic human pose sequences [53]. This field has seen substantial advancements with the rise of deep generative models [22, 33, 9, 16]. Previous works have explored text-conditioned motion generation that transform narrative descriptions into coherent pose sequences [38, 37, 25, 31, 28], audio-conditioned methods that synchronize movements with rhythmic cues [17, 35, 40], and scene-conditioned generation that integrates environmental contexts to produce contextually appropriate motions [5, 42, 1]. Despite significant progress, current mainstream data-driven kinematic motion generation methods sometimes produce unnatural motions that are jittery, distorted, or violate physiological and physical constraints. These issues could be attributed to the inherent uncertainty of the task, limitations of supervision signals, and dataset noises. Furthermore, evaluating generated human motions presents additional challenges. Conventional metrics such as error and FID fall short in capturing the nuanced details essential for producing lifelike and visually appealing movements [40, 41]. These measures can overlook critical aspects like the fluidity and biomechanical plausibility that are fundamental to human perceptual judgments. Given these challenges, it is imperative to develop metrics that are more closely aligned with human perception to more accurately evaluate and enhance the motion generation results.
@@ 2.2 Human Perception Modeling
Pioneer work [51] collect human perceptual similarity dataset and propose to utilize distance in deep features as perceptual metrics. Some works [30, 2, 50, 14, 6] in language models to explore aligning model performance with human intent by first training a reward model, then performing reinforcement learning with the reward model. Recent works  [49, 43, 24] also explore utilizing human feedback to improve visual generation results. For example, ImageReward [44] propose a reward feedback learning method (ReFL) to to align text-to-image generative models with human judgements. In human motion generation, however, few studies have explored modeling human feedbacks, even though the generated motion quality is highly relevant to human perceptions. One recent work, MoBERT [41], constructs a dataset of human ratings for generated motions. Our work differs from MoBERT in that we collect real human data on a scale tens of times larger (52.6K vs 1.4K) and use comparisons instead of ratings, which is more robust. We design the critic model to learn ratings from these comparisons automatically. Additionally, our approach could not only evaluate motion quality but also effectively improve motion generation results.
@@ 3.1 Motion Data Collection
We first collect generated human motion sequence pairs for subsequent perceptual evaluation. We utilize state-of-the-art diffusion-based motion generation method MDM [38] and FLAME [21] to generate human motion sequences parameterized by SMPL [27]. For MDM [38], we utilize the action-to-motion model trained on HumanAct12 [13] and UESTC [19] respectively. For FLAME [21], we utilize the text-to-motion model trained on HumanML3D [12]. For each group of $4$ motion sequences to be annotated, we use the same condition (text prompt or action labels) while sampling different random noises. This makes the motions similar in content while still having distinguishable differences, thereby making it easier to annotate the choices.
@@ 3.2.1 Question Design
Our perceptual evaluation is designed in the form of multiple-choice questions as selection is generally easier and more robust than directly rating [20, 36, 41]. Given a group of four motion sequence options, we instruct the annotators to select the best candidate that is most natural, visually pleasing, and free of artifacts. Specifically, we summarize the typical failure modes of the generated motions (e.g., jittering, foot skating, limb distortion, penetration, etc.) and explicitly require the annotators to exclude these options. We provide detailed guidance with task descriptions and representative video examples to better communicate the goal to the annotators. The full guidance is presented in the supplementary materials. While the optimal choice can be decided unambiguously in most cases, there are situations where the decision can be challenging. Therefore, we add two additional options, “all good” and “all bad”, so that the annotator is not required to pick one of the motions in these cases, thereby improving overall annotation quality. Results indicate that these cases account for a small portion of the total data. We exclude these cases from our subsequent experiments. In total, we set six options for each entry: four motion candidates plus “all good” and “all bad”.
@@ 3.2.2 Protocols
To ensure the quality of perceptual evaluation results, our annotation process consists of annotator training, annotation, and quality control. We recruit 10 annotators to perform the perceptual evaluation. Before the evaluation begins, we provide annotation guidelines to help the annotators understand the task and maintain consistent criteria. The annotators must pass a pilot test before starting the formal annotation to ensure they correctly understand the annotation requirements. Additionally, we conduct a perceptual consensus experiment to assess whether the annotation pipeline is suitable for our dataset, as discussed in Section 3.3. Finally, we implement a quality control process where the annotated data is reviewed by an expert quality inspector. During the annotation process, we continuously monitor the quality of each batch of data. For each batch, we randomly sample 10% of the data for quality inspection. The consistency between the sampled data and the expert’s annotations must exceed 90%; otherwise, the entire batch will be re-annotated.
((A))
((B))
((C))
@@ 3.3 Analysis
In total, we collect annotations for 18260 multiple-choice questions covering 73040 unique motions, significantly surpassing previous work [41] (1400 motions). We further investigate the following two questions:
@@ 4.1 Problem Formulation
We formulate the problem as follows: given an input human motion sequence $\mathbf{x}$, we assume there is an implicit human perception model $\mathcal{H}$ that rates the motion quality $\mathcal{H}(\mathbf{x})$, where a higher rate indicates better quality. We aim to build a computational critic model $\mathcal{C}$ that best aligns with $\mathcal{H}$. Since $\mathcal{H}$ is not explicitly available, we take a data-driven approach. We obtain the human perceptual evaluation dataset $\mathcal{D}$ containing multiple pairs of samples $(\mathbf{x}^{(i)},\mathbf{x}^{(j)})$. Our training objective is to train the model $\mathcal{C}$ using the dataset $\mathcal{D}$ so that it approximates the human perception model $\mathcal{H}$ as closely as possible. Specifically, we want the model prediction $\mathcal{C}(\mathbf{x}^{(i)})>\mathcal{C}(\mathbf{x}^{(j)})$ if and only if $\mathcal{H}(\mathbf{x}^{(i)})>\mathcal{H}(\mathbf{x}^{(j)})$. Based on the Bradley-Terry model [4, 18], the overall training objective could be written as maximizing the joint probabilities that the model $\mathcal{C}$ makes judgments consistent with $\mathcal{H}$ for each pair of samples in the dataset $\mathcal{D}$:
@@ 4.3 Motion Generation with Critic Model Supervision
Additionally, we explore to utilize the learned human perceptual prior of $\mathcal{C}$ not only for evaluating generated motions, but also improving them. We demonstrate that our motion critic model could be integrated into state-of-the-art diffusion-based motion generation approaches with ease by using MDM [38] as an example. The forward diffusion is modeled as a Markov noising process $\{\mathbf{x}_{t}\}_{t=0}^{T}$ where $\mathbf{x}_{0}$ is drawn from the data distribution, and
@@
1:Dataset: Action-label set $\mathcal{Y}=\left\{y_{1},y_{2},...,y_{n}\right\}$
2:Pre-training Dataset: Action-motion pairs dataset $\widetilde{\mathcal{D}}=\{(\textrm{label}_{1},\textrm{mot}_{1}),...,(\textrm{label}_{n},\textrm{mot}_{n})\}$
@@
3:Input: MDM model $\mathcal{M}_{\theta_{0}}$ with pre-trained weights $\mathcal{\theta}_{0}$, Critic model $\mathcal{C}$, MDM loss function $\psi$, Critic-to-loss map function $\phi$, Critic re-weight scale $\lambda$, KL loss re-weight scale $\mu$
4:Initialization: The number of noise scheduler time steps $T$, and time step range for fine-tuning $[T_{1},T_{2}]$
5:for  $(\textrm{label}_{i},\textrm{mot}_{i})\in\widetilde{\mathcal{D}}$ do
6:     $\mathcal{L}_{\text{MDM}}\leftarrow\psi_{\theta_{i}}(\textrm{label}_{i},\textrm{mot}_{i})$
@@
7:     $\theta_{i}\leftarrow\theta_{i}$triangleright Update $\textrm{MDM}_{\theta_{i}}$ with $\mathcal{L}_{\text{MDM}}$
@@
8:     $t\leftarrow\text{rand}(T_{1},T_{2})$triangleright Pick a random time step $t\in[T_{1},T_{2}]$
9:     $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$triangleright Sample noise
10:     for $j=T,...,t+1$ do
11:         no grad: $\mathbf{x}_{j-1}\leftarrow\mathcal{M}_{\theta_{i}}\{\mathbf{x}_{j}\}$
12:     end for
@@
13:     with grad: $\mathbf{\widetilde{\mathbf{x}_{0}}}^{\prime}\leftarrow\mathcal{M}_{\theta_{i}}\{\mathbf{x}_{t}\}$
14:     if $\widetilde{\mathbf{x}_{0}}^{\prime}$ is not None then
@@
15:         $\mathcal{L}_{KL}\leftarrow\mu\text{KL}(\widetilde{\mathbf{x}_{0}}^{\prime},\mathbf{x}_{0}\textquoteright$ ’ $\triangleright$ KL loss with previous $\mathbf{x}_{0}\textquoteright$ ’
16:     end if
17:     $r\sim\mathcal{N}(0,1)$triangleright Random noise scalar
@@
18:     $\mathcal{L}_{\text{Critic}}\leftarrow\lambda\phi(\mathcal{C}(\mathbf{x}_{0}\textquoteright),r)$triangleright Critic loss
@@
19:     $\theta_{i+1}\leftarrow\theta_{i}$triangleright Update $\mathcal{M}_{\theta_{i}}$ with $\mathcal{L}_{\text{Critic}}$ and $\mathcal{L}_{\text{KL}}$
@@
20:     $\widetilde{\mathbf{x}_{0}}^{\prime}\leftarrow\mathbf{x}_{0}\textquoteright$ ’ $\triangleright$ Save $\mathbf{x}_{0}\textquoteright$ ’ for next-step $\mathcal{L}_{\text{KL}}$
21:end for
@@
Given an MDM model $\mathcal{M}$ with pre-trained parameters $\theta_{0}$, we fine-tune to improve its alignment with a pre-trained critic model $\mathcal{C}$. We develop a lightweight perceptual-aligned fine-tuning approach based on ReFL [44]. Notably, in order to utilize the critic model in a plug-and-play manner, we keep the MDM training step and objective $\mathcal{L}_{\text{MDM}}$ unchanged. Instead, we simply add one optimization step with critic model supervision in each training iteration as shown in Figure 1.
@@
Specifically, we sample a Gaussian noise $\mathbf{x}_{T}$ and inference until $\mathbf{x}_{t}$ where $t\in[T_{1},T_{2}]$ is randomly selected in later denoising steps. Then, a single-step denoising is performed to predict $\mathbf{x}_{0}\textquoteright$ ’ from $\mathbf{x}_{t}$. Based on the predicted motion $\mathbf{x}_{0}\textquoteright$ ’, we compute the critic score $s=\mathcal{C}(\mathbf{x}_{0}\textquoteright)$ ’ ), which is used to compute the motion critic loss:
@@
where $\phi(s)=-\sigma(\tau-s))$ is a critic-to-loss mapping function, $\tau$ is a constant for shifting the critic value, $\sigma$ is the sigmoid function. We further introduce a Kullback-Leibler (KL) divergence regularization to prevent $\mathcal{M}$ from moving substantially away from the conditional motion generation task:
@@
The overall fine-tuning loss is given by
@@
where $\lambda$ and $\mu$ are constants for loss balancing. The detailed algorithm workflow is shown in Algorithm 1.

### v2

@@ 1 Introduction
Human motion generation is an important emerging task (Zhu et al., 2024) with wide-ranging applications, including augmented and virtual reality (AR/VR) (Yin et al., 2022; Yang et al., 2020), human-robot interaction (Nishimura et al., 2020; Gulletta et al., 2020), and digital humans (Kucherenko et al., 2019; Yi et al., 2023). Achieving high realism in generated human motions is crucial, necessitating naturalness, smoothness, and plausibility. However, current generation methods still fall short of these goals, often producing subpar results. Meanwhile, designing appropriate evaluation metrics that accurately reflect these qualities remains a significant challenge. This complexity stems from the highly non-linear and articulated nature of human motion, which must adhere to physical and bio-mechanical constraints while also avoiding visual artifacts. Effective metrics would not only facilitate the objective comparison of generated results but also have the potential to enhance generation models by addressing their shortcomings.
@@
Existing evaluation metrics typically rely on error with pairing ground truth (GT) motion, simple heuristics, or on distribution distance with real motion manifold. The error-based metrics cannot fully reflect the performance because GT is only one reasonable possibility. The heuristics fall short in comprehensively representing motion quality. For instance, foot-ground contact metrics (Rempe et al., 2021; Tseng et al., 2023) fail to penalize twisting arm motions that violate bio-mechanical constraints. It is also infeasible to manually define all the human motion rules in a handcrafted manner. Meanwhile, distribution distance metrics like Fréchet Inception Distance (FID) (Heusel et al., 2017) do not operate on an instance level but rather assess overall distribution similarity. Consequently, they cannot identify implausible motions or provide direct supervision signals to guide the generation of higher-quality motions. Some studies (Tseng et al., 2023; Voas et al., 2023) also indicate that FID correlates poorly with user studies due to the misalignment between its distance measurement and human perception of motion quality. Consequently, existing automatic evaluation metrics cannot effectively reflect or replace subjective user studies, hindering objective evaluation and comparison.
@@
In light of this, we advocate the need for automatic evaluation aligned with human perceptions. Firstly, humans are the primary audience and interaction partners for motion generation, making their perception crucial for evaluating motion quality. Secondly, the human brain possesses specialized neural mechanisms for processing biological motion (Blakemore & Decety, 2001; Grossman et al., 2000) and is sensitive to even slightly unnatural motions (Troje, 2002; Shimada & Oki, 2012). Therefore, we explore the possibility of directly learning perceptual evaluations from humans using a data-driven approach. This method could bridge the gap between objective metrics and subjective human judgments, providing a more accurate assessment of motion quality.
@@
First, we carefully curate a human perceptual evaluation dataset named MotionPercept, which contains $52563$ pairs of human preference annotations on generated motions. Next, we train a human motion critic model, MotionCritic, that learns motion quality ratings from the collected dataset. Our critic model significantly outperforms previous metrics in terms of alignment with human perceptions. Notably, it generalizes well across different data distributions. In addition to motion evaluation, we further propose to utilize the critic model as a direct supervision signal. We demonstrate that MotionCritic can be seamlessly integrated into the generation training pipeline, effectively improving motion generation quality by increasing alignment with human perceptions with few steps of finetuning.
@@ 2.1 Human Motion Generation
Human motion generation is a pivotal task in computer vision, computer graphics, and artificial intelligence, aiming to produce natural and realistic human pose sequences (Zhu et al., 2024). This field has seen substantial advancements with the rise of deep generative models (Kingma & Welling, 2014; Rezende & Mohamed, 2015; Goodfellow et al., 2014; Ho et al., 2020). Previous works have explored text-conditioned motion generation that transform narrative descriptions into coherent pose sequences (Tevet et al., 2023; 2022; Li et al., 2017; Petrovich et al., 2022; Lucas* et al., 2022), audio-conditioned methods that synchronize movements with rhythmic cues (Huang et al., 2021; Siyao et al., 2022; Tseng et al., 2023), and scene-conditioned generation that integrates environmental contexts to produce contextually appropriate motions (Corona et al., 2020; Wang et al., 2021; Araújo et al., 2023). Despite significant progress, current mainstream data-driven kinematic motion generation methods sometimes produce unnatural motions that are jittery, distorted, or violate physiological and physical constraints. These issues could be attributed to the inherent uncertainty of the task, limitations of supervision signals, and dataset noises. Furthermore, evaluating generated human motions presents additional challenges. Traditional metrics like error and FID fail to capture key aspects such as fluidity and biomechanical plausibility. Some works incorporate handcrafted physical priors (Tseng et al., 2023; Rempe et al., 2021), while others propose improved distance metrics for human poses or motions (Gopalakrishnan et al., 2019; Tanke et al., 2021; Tiwari et al., 2022). However, they still face limitations in fully and accurately reflecting human evaluations of motion quality. These challenges highlight the need for metrics that better align with human perception to more effectively evaluate and improve motion generation results.
@@ 2.2 Human Perception Modeling
Pioneer work (Zhang et al., 2018) collect human perceptual similarity dataset and propose to utilize distance in deep features as perceptual metrics. Some works (Ouyang et al., 2022; Bai et al., 2022; Yuan et al., 2023b; Hejna & Sadigh, 2024; Dong et al., 2023) in language models to explore aligning model performance with human intent by first training a reward model, then performing reinforcement learning with the reward model. Recent works  (Yuan et al., 2023a; Wu et al., 2023; Lee et al., 2023) also explore utilizing human feedback to improve visual generation results. For example, ImageReward (Xu et al., 2024) propose a reward feedback learning method (ReFL) to to align text-to-image generative models with human judgements. In human motion generation, however, few studies have explored modeling human feedbacks, even though the generated motion quality is highly relevant to human perceptions. HuTuMotion (Han et al., 2024) proposes to leverage few-shot human feedback with a set of representative texts. One recent work, MoBERT (Voas et al., 2023), constructs a dataset of human ratings for generated motions. Our work differs from MoBERT in that we collect real human data on a scale tens of times larger (52.6K vs 1.4K) and use comparisons instead of ratings, which is more robust. We design the critic model to learn ratings from these comparisons automatically. Additionally, our approach could not only evaluate motion quality but also effectively improve motion generation results.
@@ 3.1 Motion Data Collection
We first collect generated human motion sequence pairs for subsequent perceptual evaluation. We utilize state-of-the-art diffusion-based motion generation method MDM (Tevet et al., 2023) and FLAME (Kim et al., 2023) to generate human motion sequences parameterized by SMPL (Loper et al., 2015). For MDM (Tevet et al., 2023), we utilize the action-to-motion model trained on HumanAct12 (Guo et al., 2020) and UESTC (Ji et al., 2018) respectively. For FLAME (Kim et al., 2023), we utilize the text-to-motion model trained on HumanML3D (Guo et al., 2022). For each group of $4$ motion sequences to be annotated, we use the same condition (text prompt or action labels) while sampling different random noises, with a same length of 60 frames, 24 fps. This makes the motions similar in content while still having distinguishable differences, thereby making it easier to annotate the choices.
@@ 3.2.1 Question Design
Our perceptual evaluation is designed in the form of multiple-choice questions as selection is generally easier and more robust than directly rating (Kendall, 1948; Stewart et al., 2005; Voas et al., 2023). Given a group of four motion sequence options, we instruct the annotators to select the best candidate that is most natural, visually pleasing, and free of artifacts. Specifically, we summarize the typical failure modes of the generated motions (e.g., jittering, foot skating, limb distortion, penetration, etc.) and explicitly require the annotators to exclude these options. We provide detailed guidance with task descriptions and representative video examples to better communicate the goal to the annotators. The full guidance is presented in  Figure 8.
While the optimal choice can be decided unambiguously in most cases, there are situations where the decision can be challenging. Therefore, we add two additional options, “all good” and “all bad”, so that the annotator is not required to pick one of the motions in these cases, thereby improving overall annotation quality. Results indicate that these cases account for a small portion of the total data. We exclude these cases from our subsequent experiments. In total, we set six options for each entry: four motion candidates plus “all good” and “all bad”.
@@ 3.2.2 Protocols
To ensure the quality of perceptual evaluation results, our annotation process consists of annotator training, annotation, and quality control. We recruit 10 annotators to perform the perceptual evaluation. Before the evaluation begins, we provide annotation guidelines to help the annotators understand the task and maintain consistent criteria. The annotators must pass a pilot test before starting the formal annotation to ensure they correctly understand the annotation requirements. Additionally, we conduct a perceptual consensus experiment to assess whether the annotation pipeline is suitable for our dataset, as discussed in Section 3.3. Finally, we implement a quality control process where the annotated data is reviewed by an expert quality inspector. During the annotation process, we continuously monitor the quality of each batch of data. For each batch, we randomly sample 10% of the data for quality inspection. The consistency between the sampled data and the expert’s annotations must exceed 90%; otherwise, the entire batch will be re-annotated. Complete protocols are detailed in  Appendix A.
((A)) Number of Annotators Selecting the Most Preferred Option
((B)) Number of Options Selected Per Question
((C)) Pairwise Agreement Ratio Among Annotators
@@ 3.3 Analysis
In total, we collect annotations for 18260 multiple-choice questions covering 73040 unique motions, significantly surpassing previous work (Voas et al., 2023) (1400 motions). We further investigate the following two questions:
@@ 4.1 Problem Formulation
We formulate the problem as follows: given an input human motion sequence $\mathbf{x}$, we assume there is an implicit human perception model $\mathcal{H}$ that rates the motion quality $\mathcal{H}(\mathbf{x})$, where a higher rate indicates better quality. We aim to build a computational critic model $\mathcal{C}$ that best aligns with $\mathcal{H}$. Since $\mathcal{H}$ is not explicitly available, we take a data-driven approach. We obtain the human perceptual evaluation dataset $\mathcal{D}$ containing multiple pairs of samples $(\mathbf{x}^{(i)},\mathbf{x}^{(j)})$. Our training objective is to train the model $\mathcal{C}$ using the dataset $\mathcal{D}$ so that it approximates the human perception model $\mathcal{H}$ as closely as possible. Specifically, we want the model prediction $\mathcal{C}(\mathbf{x}^{(i)})>\mathcal{C}(\mathbf{x}^{(j)})$ if and only if $\mathcal{H}(\mathbf{x}^{(i)})>\mathcal{H}(\mathbf{x}^{(j)})$. Based on the Bradley-Terry model (Bradley & Terry, 1952; Hunter, 2004), the overall training objective could be written as maximizing the joint probabilities that the model $\mathcal{C}$ makes judgments consistent with $\mathcal{H}$ for each pair of samples in the dataset $\mathcal{D}$:
@@ 4.3 Motion Generation with Critic Model Supervision
Additionally, we explore to utilize the learned human perceptual prior of $\mathcal{C}$ not only for evaluating generated motions, but also improving them. We demonstrate that our motion critic model could be integrated into state-of-the-art diffusion-based motion generation approaches with ease by using MDM (Tevet et al., 2023) as an example. The forward diffusion is modeled as a Markov noising process $\{\mathbf{x}_{t}\}_{t=0}^{T}$ where $\mathbf{x}_{0}$ is drawn from the data distribution, and
@@
1:Dataset: Action-label set $\widetilde{\mathcal{D}}=\{(\textrm{label}_{i},\textrm{mot}_{i})\}$
@@
2:Input: MDM model $\mathcal{M}_{\theta_{0}}$, Critic model $\mathcal{C}$, Critic-to-loss map function $\phi$, Critic loss scale $\lambda$, KL loss scale $\mu$
3:for  $(\textrm{label}_{i},\textrm{mot}_{i})\in\widetilde{\mathcal{D}}$ do
@@
4:     $\theta_{i}\leftarrow\theta_{i}$triangleright Update $\textrm{MDM}_{\theta_{i}}$ with $\mathcal{L}_{\text{MDM}}$
@@
5:     $t\leftarrow\text{rand}(T_{1},T_{2})$triangleright Pick a random time step $t\in[T_{1},T_{2}]$
6:     $\mathbf{x}_{T}\sim\mathcal{N}(0,I)$triangleright Sample noise
7:     for $j=T,...,t+1$ do
8:         no grad: $\mathbf{x}_{j-1}\leftarrow\mathcal{M}_{\theta_{i}}\{\mathbf{x}_{j}\}$
9:     end for
@@
10:     with grad: $\mathbf{\widetilde{\mathbf{x}_{0}}}^{\prime}\leftarrow\mathcal{M}_{\theta_{i}}\{\mathbf{x}_{t}\}$
11:     if $\widetilde{\mathbf{x}_{0}}^{\prime}$ is not None then
@@
12:         $\mathcal{L}_{\text{KL}}\leftarrow\mu\text{KL}(\widetilde{\mathbf{x}_{0}}^{\prime},\mathbf{x}_{0}\textquoteright$ ’ $\triangleright$ KL loss with previous $\mathbf{x}_{0}\textquoteright$ ’
13:     end if
@@
14:     $\mathcal{L}_{\text{Critic}}\leftarrow\lambda\phi(\mathcal{C}(\mathbf{x}_{0}\textquoteright))$ ’ ) ) $\triangleright$ Critic loss
@@
15:     $\theta_{i+1}\leftarrow\theta_{i}$triangleright Update $\mathcal{M}_{\theta_{i}}$ with $\mathcal{L}_{\text{Critic}}$ and $\mathcal{L}_{\text{KL}}$
@@
16:     $\widetilde{\mathbf{x}_{0}}^{\prime}\leftarrow\mathbf{x}_{0}\textquoteright$ ’ $\triangleright$ Save $\mathbf{x}_{0}\textquoteright$ ’ for next-step $\mathcal{L}_{\text{KL}}$
17:end for
@@
Given an MDM model $\mathcal{M}$ with pre-trained parameters $\theta_{0}$, we fine-tune to improve its alignment with a pre-trained critic model $\mathcal{C}$. We develop a lightweight perceptual-aligned fine-tuning approach based on ReFL (Xu et al., 2024). Notably, in order to utilize the critic model in a plug-and-play manner, we keep the MDM training step and objective $\mathcal{L}_{\text{MDM}}$ unchanged. Of each fine-tuning iteration, the original MDM loss is computed using text labels and ground-truth motions as the initial step. This loss is first utilized to update the pretrained model weights accordingly. After that, we simply add another optimization step with critic model supervision. Structure of this optimization step is shown in Figure 3, and detailed as follows:
@@
We begin with sampling a Gaussian noise $\mathbf{x}_{T}$ and performing gradient-free denoising steps until $\mathbf{x}_{t}$, where $t\in[T_{1},T_{2}]$ is randomly selected in denoising steps. Gradient interception range $[T_{1},T_{2}]$ are hyperparameters, with selection principles outlined in Section C.1. After reaching step $t$, a single-step denoising with gradient is performed directly to predict $\mathbf{x}_{0}\textquoteright$ ’ from $\mathbf{x}_{t}$. Upon the predicted motion $\mathbf{x}_{0}\textquoteright$ ’, we compute its critic score $s=\mathcal{C}(\mathbf{x}_{0}\textquoteright)$ ’ ), and use this score output to compute the motion critic loss. We formulate our critic loss as follows:
@@
where $\phi(s)=-\sigma(\tau-s))$ is a critic-to-loss mapping function, $\tau$ being a constant threshold for shifting the critic value, and $\sigma$ being sigmoid function.
We further introduce a Kullback-Leibler (KL) divergence regularization to prevent $\mathcal{M}$ from moving substantially away from the conditional motion generation task. We formulate our KL loss as follows:
@@
where $\widetilde{\mathbf{x}_{0}^{\prime}}$ being $\mathbf{x}_{0}^{\prime}$ of the previous iteration. Overall, our fine-tuning loss is given by
@@
where $\lambda$ and $\mu$ are re-scaling weights for loss balancing. Detailed algorithm workflow is shown in Algorithm 1.